OpenStudy (anonymous):

Deb and Sam go to the mall, each have $200 to spend. Deb spent $150 and Sam spent $160. What percent of their total money did they spend?

OpenStudy (anonymous):

150/200 gives you what ? It gives you 0.75. Now multiply that by 100 and you will have your percent. 0.75 x 100 = 75% now you try the other one... 160/200 = try it and see what you get.

OpenStudy (anonymous):

Thanks So Much That Was Correct And Im About To Solve It And Get Back To You :)))

OpenStudy (anonymous):

I read the question wrong...sorry about that total money = 400 150 + 160 = 310 310/400 = 0.775 0.775 x 100 = 77.5%